OpenStudy (anonymous):

the access code for a car's security system consists of four digits. the first digit cannot be 1 and the last digit must be even or zero. how many different codes are available?

OpenStudy (shane_b):

1st value: 1 in 9 2nd value: 1 in 10 3rd value 1 in 10 4th value: 1 in 5 \[9*10*10*5=?\]

OpenStudy (anonymous):

thanks so much shane. that looks right!

OpenStudy (shane_b):

yw :)

OpenStudy (anonymous):

here's one more i can't seem to figure out, if you're willing. A certain lottery has 30 numbers. In how many different ways can 6 of the numbers be selected? Assume that order of selection is not important.

OpenStudy (shane_b):

Can each number be selected more than once?

OpenStudy (anonymous):

the question does not say. i typed it word for word...

OpenStudy (shane_b):

My assumption is that each number can only be selected once...since it's a lottery

OpenStudy (anonymous):

i don't think they can be selected more than once because it said 6 numbers, so think your assumption is correct.

OpenStudy (shane_b):

I'm no probability expert but I believe it would simply be: \[30*29*28*27*26*25=?\]

OpenStudy (anonymous):

that looks to be 427,518,000. the answer is supposed to be 593,775. can't figure out how they got that answer.

OpenStudy (shane_b):

I guess since the order doesn't matter it would be: \[\frac{n!}{r!(n-r)!}=\frac{30!}{6!(30-6)!}\]

OpenStudy (shane_b):

Which = 593,775

OpenStudy (anonymous):

but how do you figure that exactly. i have a calculator with the x! function, but the number i get does not resemble this, so i must be doing something wrong.

OpenStudy (shane_b):

I'm not sure how to do it on my calculator either (TI-84). I just used Microsoft Math to check it. It's worth the download and it's free. http://www.microsoft.com/en-us/download/details.aspx?id=15702

OpenStudy (anonymous):

ok i'll try it - thanks a million.

OpenStudy (shane_b):

np :)

OpenStudy (anonymous):

ok. i downloaded the calculator. handy little thing. but i still can't figure out what to input to get this answer. i typed in {30!}{6!(30-6)!} i got back some long decimal answer. any suggestions?

OpenStudy (shane_b):

You're missing the division sign :)